# Score-based Generative Diffusion Models for Social Recommendations

### v1

@@ II-D Curriculum-Driven Training of the Social Denoising Model
Since the optimal social representation that maximizes consistency with collaborative signals is difficult to determine, training the SGM for social graph denoising presents significant challenges due to the lack of ground truth and supervision signals. For example, in image restoration tasks, SGMs use deteriorated images as input and compute the loss function between the ground-truth image and the denoised image. This straightforward training strategy is inapplicable in the context of social denoising, as the ground-truth optimal social representation is not accessible. To overcome this limitation, we employ a curriculum learning strategy that enables the SGM to perceive social representations across social graphs with varying levels of sparsity. This approach potentially encompasses the optimal social graph, which the SGM model can identify automatically with collaborative signals.
@@
  • In the initial stage, the model retains the percentage $\rho$ of social connections that exhibit the most similar preference patterns for $N$ epochs. Given the collaborative representations of user $i$ and $j$ from $RecEnc(\cdot)$ as $(\mathbf{e}_{R}^{(L)})_{u_{i}}$ and $(\mathbf{e}_{R}^{(L)})_{u_{j}}$, we leverage the inner product to measure the similarity ${sim}=(\mathbf{e}_{R}^{(L)})_{u_{i}}\cdot{(\mathbf{e}_{R}^{(L)})^{T}_{u_{j}}}$  [18].
@@
$\min_{\boldsymbol{\theta}}\mathbb{E}_{t}\left[\lambda(t)\mathbb{E}_{\mathbf{\bar{x}}_{0}}\mathbb{E}_{\mathbf{\bar{x}}_{t}|\mathbf{\bar{x}}_{0},\mathbf{c}}\left\lVert{\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{c})\!-\!{\nabla}_{\mathbf{x}_{t}}{\log p_{t}(\mathbf{\bar{x}}_{t}|\mathbf{\bar{x}}_{0})}}\right\rVert^{2}\right],$ (11)
where $\bar{\mathbf{x}}_{0}=(\mathbf{\bar{e}}_{S}^{(L)})_{u}\sim\bar{\mathbf{E}}_{S}^{(L)}$ denotes the social embedding of sampled user $u$, $\mathbf{\bar{x}_{t}}\sim p_{0t}(\mathbf{\bar{x}_{t}}|\mathbf{\bar{x}_{0}})$ represents the distribution of data at the intermediate diffusion step $t$, $\mathbf{c}$ representes the corresponding collaborative representation as disentangled condition. Note that the SGSR does not designate a specific sparsity level as the optimal representation; instead, it is determined by the recommendation signal through joint training strategy as described in  2.
II-E Score-based Denoising Model Optimization
@@
     if epoch == 0 or epoch %$\%(M+N)$ == 0 then
@@
     else if epoch %$\%N$ == 0: then
@@
ener process and exposure bias in the interactive sampling process may introduce undesired noise to the generated social representation  [44, 45]. To mitigate these biases and effectively leverage the learned social representation distribution, the SGSR incorporates contrastive learning techniques, known for their high efficiency in knowledge transfer  [46, 16]. The SGSR enhances the RS with social information by maximizing the consistency of user representations across the social and collaborative domains.
@@
$\mathcal{L}^{cl}=\sum_{u\in\mathcal{U}}-\log\frac{\varsigma(\mathbf{\hat{e}}_{u},\mathbf{p}_{u})}{\varsigma(\mathbf{\hat{e}}_{u},\mathbf{p}_{u})+\sum\limits_{k\neq j}\varsigma(\mathbf{\hat{e}}_{u},\mathbf{p}_{u}))}$ (12)
where $\mathbf{\hat{e}}_{u}=MLP\left((\mathbf{\hat{e}}_{S}^{(L)})_{u}\right)$ and $\mathbf{p}_{u}=MLP\left((\mathbf{e}_{R}^{(L)})_{u}\right)$ denotes the social representation and collaborative representation of user $u$ projected to the same semantic space by multi-layer perceptions $MLP(\cdot)$, $\varsigma(\mathbf{\hat{e}}_{u},\mathbf{p}_{u})=\exp\left({\cos(\mathbf{\hat{e}}_{u},\mathbf{p}_{u})}/\tau\right)$, and $\tau$ is the temperature parameter.
@@ III-A Experiment Settings
(2) Evaluation Metrics. To evaluate top-$K$ recommendations from implicit user feedback, we use two standard ranking metrics: Recall@$K$ (R@$K$) and Normalized Discounted Cumulative Gain (N@$K$). We set $K\in\{5,10\}$ for evaluation and rank items among all candidates to conduct a full-rank evaluation protocol. To ensure the stability of the results, each experiment is repeated 5 times, and we report the average values.
(3) Baselines. We evaluate the SGSR’s performance against the state-of-the-art baselines across several categories: vanilla social recommendation models (SocialMF [49], NGCF [2], LightGCN [3], DiffNet [32], SEPT[50], MHCN [46]), diffusion recommendation models (DiffRec [51], DreamRec [52], DreamRec+), and social graph denoising methods (DESIGN [19], DSL [53], GBSR [20], SHaRe [11]). Descriptions of these baseline methods are provided as follows:
@@
  • DreamRec+: DreamRec+ combines the social information with the generation guidance embedding to adopt the DreamRec to social recommendation
  • SocialMF  [49]: SocialMF propagates the social information into the matrix factorization model
@@
  • GBSR [20]): GBSR introduces a theoretically-inspired framework to maximize the mutual information between the denoised social graph and the interaction matrix
  • SHaRe [11]: SHaRe designs the graph rewiring technique to discretely edit the social graphs based on the cosine similarity between user’s collaborative representations.
[…]
@@ III-B Overall Performance
We evaluate the recommendation performance of all baselines and our proposed approach SGSR. Table II summarizes the Recall and NDCG score across the Ciao, Epinions, and Dianping datasets. Our main findings include:
@@
  • In most cases, social denoising methods surpass the GNN-based social recommendation that directly utilizes the raw social graph. This observation underscores the importance of addressing the challenge of low social homophily. However, the performance of these social recommendation methods is sub-optimal compared with the proposed SGSR, likely due to their reliance on simple, handcrafted similarity metrics and the restricted scope of rewired edges.
@@
  • Diffusion-based methods underperform most vanilla social recommendation models. This observation indicates that existing diffusion models for recommendation cannot be simply adapted to the social recommendation task, necessitating the need to address the unique challenges in developing a diffusion-based recommender model for social recommendations.
Figure 4: Comparision of SGSR with its variants on three datasets across metric Recall@10 and NDCG@10.
@@ III-C Ablation Study
TABLE III: Ablation Study on Generative Models
@@
TABLE IV: Effect of Different Encoders
@@
Ablation Study on Generative Models. We further conduct comparison experiments with alternatives of generative models including VAE, DDPM  [28], NCSN  [29], and different variants of SGMs. For VAE, DDPM, and NCSN, we used their official code to ensure a fair comparison. For the SGM variants, the implementation details are provided in Section  III-D, achieving different instantiations of SGM with various drift and diffusion coefficients. The results are summarized in the Table  III.
@@
Ablation Study on Graph Encoders. Graph encoders are employed to extract collaborative and social information from graph-structured data. We evaluate SGSR with various recommendation backbones to assess the impact of different encoders, following the designs of previous works. The results, presented in Table  IV, show that the proposed method achieves comparable performance with mainstream encoders. SGSR is adaptable to different encoders, even though each encoder produces representations of varying quality. This can be attributed to SGM, which enhances the denoised social representation with the learned data distribution.
@@ III-D Parameter Sensitivity Study
(1) Diffusion Steps. We tune the scale of the continuous time $T$, known as diffusion steps, on two well-established formulations: VP-SDE and VE-SDE. The results are presented in Fig. 5. For all the datasets, we observe optimal performance at 5 diffusion steps. Performance degrades as the number of time steps increases, potentially due to the introduction of additional noise in longer diffusion processes. Moreover, VP-SDE outperforms VE-SDE in most cases.
@@
(3) Maintain Rate in Curriculum Learning Mechanism. We varied the maintain rate $\rho$ and $\hat{\rho}$ within the set $\{0.7,0.75,0.8,0.85,0.9\}$. The results are presented in Fig.  7. SGSR performance peaks when we remove $20\%$ and $25\%$ of social relations for the Dianping dataset during the two-stage curriculum learning. This underscores the necessity of social graph denoising, as including the entire raw social graph would degrade recommendation performance.

### v2

@@ II-D Curriculum-Driven Training of the Social Denoising Model
Since the optimal social representation that maximizes consistency with collaborative signals is difficult to determine, training the SGM for social graph denoising presents significant challenges due to the lack of ground truth and supervision signals. For example, in image restoration tasks, SGMs use deteriorated images as input and compute the loss function between the ground-truth image and the denoised image. This straightforward training strategy is inapplicable in the context of social denoising, as the ground-truth optimal social representation is not accessible. To overcome this limitation, we employ a curriculum learning strategy that enables the SGM to perceive social representations across social graphs with varying levels of sparsity. This approach potentially encompasses the optimal social representations, which the SGM model can identify automatically with collaborative signals.
@@
  • In the initial stage, the model retains the percentage $\rho$ of social connections that exhibit the most similar preference patterns for $N$ epochs. Given the collaborative representations of user $i$ and $j$ from $RecEnc(\cdot)$ as $(\mathbf{e}_{R}^{(L)})_{u_{i}}$ and $(\mathbf{e}_{R}^{(L)})_{u_{j}}$, we leverage the inner product to measure the similarity $\mathrm{sim}=(\mathbf{e}_{R}^{(L)})_{u_{i}}\cdot{(\mathbf{e}_{R}^{(L)})^{T}_{u_{j}}}$.
@@
$\hskip-5.69054pt\min_{\boldsymbol{\theta}}\mathbb{E}_{t}\left[\lambda(t)\mathbb{E}_{\mathbf{\bar{x}}_{0}}\mathbb{E}_{\mathbf{\bar{x}}_{t}|\mathbf{\bar{x}}_{0},\mathbf{c}}\left\lVert{\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{c})\!-\!{\nabla}_{\mathbf{x}_{t}}{\log p_{0t}(\mathbf{\bar{x}}_{t}|\mathbf{\bar{x}}_{0})}}\right\rVert^{2}\right],$ (11)
where $\bar{\mathbf{x}}_{0}=(\mathbf{\bar{e}}_{S}^{(L)})_{u}\sim\bar{\mathbf{E}}_{S}^{(L)}$ denotes the social embedding of sampled user $u$, $\mathbf{\bar{x}_{t}}\sim p_{0t}(\mathbf{\bar{x}_{t}}|\mathbf{\bar{x}_{0}})$ represents the distribution of data at the intermediate diffusion step $t$, $\mathbf{c}$ representes the corresponding collaborative representation as disentangled condition. Note that the SGSR does not designate a specific sparsity level as the optimal representation; instead, it is determined by the recommendation signal through joint training strategy as described in  II-E.
@@
  if epoch == 0 or epoch $\%(M+N)$ == 0 then
@@
  else if epoch $\%N$ == 0: then
@@ II-E Score-based Denoising Model Optimization
ener process and exposure bias in the interactive sampling process may introduce undesired noise to the generated social representation [36]. To mitigate these biases and effectively leverage the learned social representation distribution, the SGSR incorporates contrastive learning techniques, known for their high efficiency in knowledge transfer  [37, 15]. The SGSR enhances the RS with social information by maximizing the consistency of user representations across the social and collaborative domains.
@@
$\mathcal{L}^{cl}=\sum_{u\in\mathcal{U}}-\log\frac{\varsigma(\mathbf{\hat{e}}_{u},\mathbf{p}_{u})}{\varsigma(\mathbf{\hat{e}}_{u},\mathbf{p}_{u})+\sum\limits_{v\in\mathcal{U},v\neq u}\varsigma(\mathbf{\hat{e}}_{u},\mathbf{p}_{v}))},$ (12)
where $\mathbf{\hat{e}}_{u}=\mathrm{MLP}\left((\mathbf{\hat{e}}_{S}^{(L)})_{u}\right)$, $\mathbf{p}_{u}=\mathrm{MLP}\left((\mathbf{e}_{R}^{(L)})_{u}\right)$, and $\mathbf{p}_{v}=\mathrm{MLP}\left((\mathbf{e}_{R}^{(L)})_{v}\right)$ denotes the projected social representation of user $u$ and collaborative representation of user $u$ and $v$, respectively.These representations are mapped into a shared semantic space using a multi-layer perception $\mathrm{MLP}(\cdot)$. The similarity function $\varsigma(\mathbf{\hat{e}}_{u},\mathbf{p}_{*})=\exp\left({\cos(\mathbf{\hat{e}}_{u},\mathbf{p}_{*})}/\tau\right)$ measures the representation similarity based on cosine similarity, where $\tau$ is the temperature parameter.
@@
II-F Complexity Analysis of SGSR
This section analyzes the space and time complexity of the proposed SGSR framework.
(1) Space Complexity Analysis. The space complexity analysis of SGSR focuses on two primary components: graph encoders and the reverse diffusion process. The social encoder $\text{{SoEnc}}(\cdot)$ requires $\mathcal{O}(|\mathbf{S}|+d^{2}+|U|d)$, assuming the input and output dimensions of GCN are equal, where $|\mathbf{S}|$ represents the number of social connections. Meanwhile, the collaborative encoder $\text{{RecEnc}}(\cdot)$ (e.g., LightGCN) needs $\mathcal{O}(|\mathbf{R}|+(|\mathcal{U}|+|\mathcal{V}|)d)$, with $|\mathbf{R}|$ representing the number of item interactions. Aditionally, the reverse diffusion process and the scoring network $\mathbf{s}_{\boldsymbol{\theta}}$ jointly require $\mathcal{O}(dT+{L}\cdot d^{2})$ space, where $T$ is the number of diffusion step, L{L} is the number of layers in the scoring network $\mathbf{s}_{\boldsymbol{\theta}}$, and $d$ represents the dimensionality of each layer, assumed to be consistent across layers for brevity.
(2) Time Complexity Analysis. From a theoretical perspective, the time complexity analysis of SGSR primarily consists of two components: GNN-based representation aggregation and diffusion-based denoising process. Initially, SGSR utilizes GCN to produce social embeddings. Each layer involves message passing and feature aggregation operations, yielding a time complexity of $\mathcal{O}(|\mathbf{S}|d+|\mathcal{U}|d^{2})$. The time complexity of $\text{{RecEnc}}(\cdot)$ depends on the specific GNN encoder. Taking LightGCN as an example, it requires $\mathcal{O}(|\mathbf{R}|d+(|\mathcal{U}|+|\mathcal{V}|)d)$. Subsequently, for each batch, the forward diffusion process requires $\mathcal{O}(Bd)$ calculations to perturb the social representation, while the one-step inference costs $\mathcal{O}(B\cdot{L}\cdot d^{2})$, where $B$ denotes the batch size.
@@ III-A Experiment Settings
(2) Evaluation Metrics. To evaluate top-$K$ recommendations from implicit user feedback, we use two standard ranking metrics: Recall@$K$ (R@$K$) and Normalized Discounted Cumulative Gain (N@$K$). We set $K\in\{5,10\}$ for evaluation and rank items among all candidates to conduct a full-rank evaluation protocol. To ensure statistical reliability, all experiments were repeated five times with results reported as mean values.
@@
TABLE III: Comparison of performance of social recommendation on Epinions datasets. The best performance is highlighted in bold, and the second best is underlined. %Improve denotes the relative improvement of SGSR over the strongest baseline.
Dataset Epinions
Method R@40 R@20 R@3 N@40 N@20 N@3
DiffNet 0.0906 0.0624 0.0169 0.0394 0.0327 0.0175
SEPT 0.0919 0.0657 0.0171 0.0417 0.0349 0.0202
MHCN 0.0939 0.0666 0.0178 0.0425 0.0352 0.0203
DESIGN 0.0935 0.0659 0.0175 0.0432 0.0353 0.0201
DSL 0.0926 0.0640 0.0153 0.0419 0.0347 0.0189
GBSR 0.0913 0.0656 0.0167 0.0427 0.0351 0.0197
SHaRe 0.0934 0.0673 0.0184 0.0430 0.0356 0.0195
EIISRS 0.0944 0.0661 0.0181 0.0431 0.0355 0.0198
RecDiff 0.0929 0.0669 0.0187 0.0426 0.0357 0.0200
GDSSL 0.0941 0.0671 0.0183 0.0429 0.0342 0.0196
SGSR 0.1013 0.0697 0.0205 0.0450 0.0372 0.0214 C
%Improve 7.31% 3.57% 9.63% 4.17% 4.20% 5.42%
(3) Baselines. We evaluate the SGSR’s performance against the state-of-the-art baselines across several categories: vanilla social recommendation models (SocialMF [40], NGCF [1], LightGCN [2], DiffNet [28], SEPT[41], MHCN [37]), diffusion recommendation models (DiffRec [42], DreamRec [43], DreamRec+, BSPM [44], GiffCF [45]), and social graph denoising methods (DESIGN [18], DSL [46], GBSR [19], SHaRe [10], EIISRS [47], RecDiff [48], GDSSL [49]). Descriptions of these baseline methods are provided as follows:
@@
  • DreamRec+: DreamRec+ combines the social information with the generation guidance embedding to adopt the DreamRec to social recommendation.
  • BSPM  [44]: BSPM applies score-based generative models to collaborative filtering by introducing a blurring and sharpening process on the interaction matrix.
  • GiffCF  [45]: GiffCF applies the diffusion algorithm utilizing the heat equation to the item-item similarity graph for collaborative filtering with implicit feedback.
  • SocialMF  [40]: SocialMF propagates the social information into the matrix factorization model.
@@
  • GBSR [19]): GBSR introduces a theoretically-inspired framework to maximize the mutual information between the denoised social graph and the interaction matrix.
  • SHaRe [10]: SHaRe designs the graph rewiring technique to discretely edit the social graphs based on the cosine similarity between users’ collaborative representations.
  • EIISRS  [47]: EIISRS refines user social representations by reconstructing social relations using a layerwise graph-enhanced variational autoencoder.
  • RecDiff  [48]: RecDiff introduces a social denoising framework that directly applies diffusion probabilistic models in the hidden space to refine encoded user representations.
  • GDSSL  [49]: GDSSL applies the discrete-time diffusion algorithm directly to the raw social adjacency matrix to mitigate social noise, subsequently employing SSL to align cross-domain information.
For all baselines, we adhere to their official implementations and recommended settings to ensure fair comparisons. Specifically, the SocialMF, NGCF, LightGCN, DiffNet, SEPT, and MHCN were implemented using the publicly available QRec  [^3] https://github.com/Coder-Yu/QRec library which is built on TensorFlow. The implementations of DiffRec, DreamRec, BSPM, GiffCF, Design, DSL, GBSR, and RecDiff were derived from their official code releases. EIISRS and GDSSL were replicated following the original papers. The DreamRec+ model integrates social information obtained through the GCN encoder $\text{{SoEnc}}(\cdot)$ as the guidance embedding for DDPM, utilizing a multilayer perceptron for this integration. Moreover, we implement social denoising baselines requiring a recommendation backbone model, as well as our proposed SGSR with the most widely used LightGCN backbone. For a fair comparison, we employ a grid search strategy to tune the hyperparameters of the baseline models, exploring the vicinity of the optimal values reported in their respective original papers. All methods are optimized using the Adam optimizer. For the diffusion-based recommenders, we meticulously tune both the upper and lower bounds of the noise, as well as the number of diffusion steps, to maximize performance. Take 5 steps in DiffRec as an example, the upper bound and lower bound are searched in $\{$0.5, 0.1, 0.05, 0.01 }\} and $\{$0.01, 0.001, 0.0001, 0.0001 }\}.
(4) Parameter Settings. The proposed framework is implemented with PyTorch. The embedding dimension $d$ is tuned within $\{$32, 64, 128, 256}\}. We use batch sizes of 1024 for Ciao and 2048 for Epinions and Dianping to improve training efficiency based on the scale of each dataset. For each positive instance, we sample one negative instance to compute the BPR loss. We use 3 GNN layers to encode the high-order social relations for our denoising framework. We employ the Adam optimizer with a learning rate of 0.001 to train the SGSR model. The diffusion step $T$ is evaluated across $\{3,5,10,30,50,100\}$ for both VE-SDE and VP-SDE formulations. We examine two predictors (reverse diffusion sampler, Euler-Maruyama) in combination with two correctors (Langevin MCMC, none). The maintain rate $\rho$ and $\hat{\rho}$ in the curriculum learning mechanism are tuned within $\{$0.7, 0.75, 0.8, 0.85, 0.9}\}. Additionally, the temperature parameter for SSL, $\tau$, which is set in $\{0.05,0.1,0.5\}$. The loss weights $\lambda_{1}$ and $\lambda_{2}$ are tuned in $\{0.001,0.01,0.1,0.5,1,5\}$.
Figure 4: Comparison of SGSR with its variants on three datasets across metric Recall@10 and NDCG@10.
@@ III-B Overall Performance
We evaluate the recommendation performance of all baselines and our proposed approach SGSR. Table II summarizes the Recall and NDCG score across the Ciao, Epinions, and Dianping datasets, while Table III demonstrates varying top-$K$ performance on the Epinions datasets, specifically evaluating social recommendation modes and graph denoising methods. Our main findings include:
@@
  • In most cases, social denoising methods surpass the GNN-based social recommendation that directly utilizes the raw social graph. This observation underscores the importance of addressing the challenge of low social homophily. While generative denoising methods obtain similar performance to graph editing techniques, SGSR shows improved performance over both. This performance gap is likely rooted in: (1) reliance on simple, handcrafted similarity metrics and restricted editing scope of graph editing methods, and (2) the inherent limitations of generative paradigms and their training paradigms in the context of social recommendation.
@@
  • Diffusion-based methods exhibit performance comparable to graph-based recommendation models, highlighting the promising potential of diffusion algorithms for recommendation tasks. Nevertheless, these methods still generally trail behind most vanilla social recommendation approaches. This observation indicates that existing diffusion models for recommendation cannot be simply adapted to the social recommendation task, necessitating the need to address the unique challenges in developing a diffusion-based recommender model for social recommendations.
@@ III-C Ablation Study
TABLE IV: Ablation Study on Generative Models
@@
TABLE V: Effect of Different Encoders
@@
Ablation Study on Generative Models. We further conduct comparison experiments with alternatives of generative models including VAE, DDPM  [24], NCSN  [25], and different variants of SGMs. For VAE, DDPM, and NCSN, we used their official code to ensure a fair comparison. For the SGM variants, the implementation details are provided in Section  III-D, achieving different instantiations of SGM with various drift and diffusion coefficients. The results are summarized in the Table  IV.
@@
Ablation Study on Graph Encoders. Graph encoders are employed to extract collaborative and social information from graph-structured data. We evaluate SGSR with various recommendation backbones to assess the impact of different encoders, following the designs of previous works. The results, presented in Table  V, show that the proposed method achieves comparable performance with mainstream encoders. SGSR is adaptable to different encoders, even though each encoder produces representations of varying quality. This can be attributed to SGM, which enhances the denoised social representation with the learned data distribution.
@@ III-D Parameter Sensitivity Study
(1) Diffusion Steps. We tune the scale of the continuous time $T$, known as diffusion steps, on two well-established formulations: VP-SDE and VE-SDE. The results are presented in Fig. 5. For all the datasets, we observe optimal performance at 5 diffusion steps. Performance degrades as the number of time steps increases, potentially due to the introduction of additional noise in longer diffusion processes. Moreover, VP-SDE demonstrates superior performance to VE-SDE in most cases, with the performance gap widening when the diffusion step ranges between 5-50. This can be attributed to VE-SDE’s tendency to significantly perturb the data distribution through high-variance noise  [23], potentially leading to personalized information loss and consequently increasing the difficulty of the reverse process.
@@
(3) Maintain Rate in Curriculum Learning Mechanism. We varied the maintain rate $\rho$ and $\hat{\rho}$ within the set $\{0.7,0.75,0.8,0.85,0.9\}$. The results are presented in Fig.  7. SGSR performance peaks when we remove $20\%$ and $25\%$ of social relations for the Dianping dataset during the two-stage curriculum learning. This suggests that retaining approximately $80\%$ of social connections during training enables SGSR to learn optimally collaborative-consistent representations. These results corroborate prior findings on social graph denoising  [17, 10], confirming that raw social graphs containing undesired noise connections negatively impact recommendation performance.